all chords of the curve 3x2-y2-2x+4y=0 that subtends a right angle at the origin pass through a point whose coordinates are.......???

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear shivanshi

let the equation of chord is y=mx+c

so combine equation od pair of straight line that passes through the origin and point of intersection of chord and  curve is

put value of 1 fron the chord  

1=(y-mx)/c

3x2-y2-(2x+4y)(y-mx)/c=0

(3c+2m)x2 + (4m-2)xy +(-c-4)y2 =0

these lines are perpendicular

so a+b =0

  3c +2m -c-4 =0

2=m*1 +c 

compair this with the chord y=mx +c

so it will always pass throug

 (1,2)

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very

quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin